OpenStudy (anonymous):

Please help i am so lost! Simplify the expression The root of negative nine over the quantity of three minus two i plus the quantity of one plus five i. .

OpenStudy (imstuck):

\[\frac{ \sqrt{-9} }{ 3-2i+(1+5i) }\]Is that it?

OpenStudy (anonymous):

Yah

OpenStudy (imstuck):

Except there was probably a set of parenthesis around the 3 - 2i, right?

OpenStudy (anonymous):

Yah there are parenthesis

OpenStudy (imstuck):

Ok, first deal with the numerator. You know that there is an "i" involved because you have a negative under the radical sign. What is the square root of 9?

OpenStudy (anonymous):

3

OpenStudy (imstuck):

And because of the - sign, it is 3i, right?

OpenStudy (anonymous):

Yah that makes sense

OpenStudy (imstuck):

You can rearrange the terms in the denominator to look like this:

OpenStudy (imstuck):

\[\frac{ 3i }{ 3+1-2i+5i }\]

OpenStudy (imstuck):

3+1 = 4 and 5i - 2i = 3i, so the new denominator looks like this:\[\frac{ 3i }{ 4+3i }\]

OpenStudy (imstuck):

You can't get any simpler than that.

OpenStudy (anonymous):

Wait a sec these are the answer choices

OpenStudy (imstuck):

ok, then I know what i did wrong.

OpenStudy (anonymous):

these are the other 3

OpenStudy (imstuck):

I failed to notice that I should have multiplied by the conjugate of the denominator. Let me work on that and I will post. I gotcha!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (imstuck):